OpenStudy (anonymous):

What is the solution of the following system 3x+3y=10 -9x-9y=-30

OpenStudy (anonymous):

I don't need the answer I just need step by step help

OpenStudy (kewlgeek555):

Okay, I am going to show you an example see if you can identify what I did. And this is linear equation, right?

OpenStudy (anonymous):

use elimination method

OpenStudy (anonymous):

Yes linear

OpenStudy (kewlgeek555):

2x + 3y = 1500 -2x -2x 3y = -2x + 1500\[\frac{ 3y }{ 3 }=\frac{ -2 }{ 3 }x+\frac{ 1500 }{ 3 }\](I am dividing here)^ So your answer comes to this: \[y=\frac{ -2 }{ 3 }x+500\] where -2 over 3 is the slope and 500 is the y-intercept.

OpenStudy (anonymous):

I'm so confused

OpenStudy (kewlgeek555):

So...did you see what I did? I started by eliminating the one to the left first. Added it to the other side (like you would do with a regular linear equation), but since you have to subtract it I added -2x. Since 3 was being multiplied by y, I divded every number by 3, which got me to my final answer. I am going to do yours now. Step by step. And then , hopefully you can do the very last step for the second one.

OpenStudy (anonymous):

Ok what if I did this (3x+3y=10)(-9) (-9x-9y=-30)(3)

OpenStudy (anonymous):

|dw:1380683040201:dw understand

OpenStudy (anonymous):

What

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

6x + 2y = 0 4x + y = -1